OpenStudy (anonymous):

simplify: 3(y-2)-2(y-6)

OpenStudy (anonymous):

2/9z = 5+3 2/9z = 8 2= 8(9z) 2= 72z z= 2/72

OpenStudy (anonymous):

Thank you

OpenStudy (anonymous):

your welcome

OpenStudy (anonymous):

2x-4<x-7 2x-x <-7-4 x<-11